When 2 dice 1 red and green are rolled there are 36 outcome. If both dice are fair the outcome are equally likely. Let E be the event that the sum of the numbers rolled is 7. What is the probability of 7 ?
1
Expert's answer
2013-03-12T13:42:22-0400
We have next combinations for the sum of the numbers rolled is 7:
6 red and 1 green: 6+1=7
5 red and 2 green: 5+2=7
4 red and 3 green:4+3=7
3 red and 4 green: 3+4=7
2 red and 5 green: 2+5=7
1 red and 6 green: 1+6=7
The total numberof combinations equals 6.
Therefore, theprobability that the sum of the numbers rolled is 7 equals:
P = 6/36 =1/6(=0.166667)
Answer: theprobability that the sum of the numbers rolled is 7 equals 1/6